# LEMUR: Learned Multi-Vector Retrieval

### v1

@@ 7 Discussion
Because of their fine-grained representation that captures token-level interactions, multi-vector models are more expressive, and, consequently, tend to have a higher retrieval quality, than single-vector models. However, the high latency of multi-vector retrieval engines has hindered their wider adoption. To address this issue, we introduce Lemur, a simple-yet-efficient framework that is an order of magnitude faster than earlier multi-vector similarity search methods, thus enabling efficient multi-vector retrieval at scale. Moreover, Lemur produces single-vector representations for both the query and the documents, reducing the task to single-vector similarity search. This enables easy integration with state-of-the-art single-vector ANNS libraries.
@@
Because they are learned in a supervised fashion, the single-vector embeddings generated by Lemur are significantly more accurate than the fixed-dimensional encodings (FDEs) (Jayaram et al., 2024) that have been used for single-vector reduction earlier: in particular, $1024$-dimensional Lemur embeddings yield higher recall than $10240$-dimensional FDEs (see Sec. 6.2). This enables using lower-dimensional single-vector embeddings, which decreases the memory footprint and leads to lower end-to-end latency due to faster single-vector similarity search. Standard scalar or product quantization (Jegou et al., 2011) techniques could be applied to the single-vector embeddings generated by Lemur to further decrease memory consumption and end-to-end latency.
@@
The performance of earlier multi-vector similarity search methods has been evaluated almost exclusively on ColBERTv2 (Santhanam et al., 2022b) embeddings. In this work, we conduct a performance evaluation on four other multi-vector text models and on two multi-vector visual document retrieval models (see Sec. 6.3). Our results indicate that the multi-vector model used to generate the embeddings has a larger impact on the performance of multi-vector similarity search methods than the embedded dataset. In particular, most of the earlier methods either fail to achieve acceptable recall levels at all, or have prohibitively high latency on the LFM2-ColBERT (Amini et al., 2025) and answerai-colbert-small (Clavié, 2025) embeddings. In contrast, Lemur consistently reaches the highest recall levels and has low latency regardless of the multi-vector model used to generate the embeddings. The above observations underline the importance of establishing robust benchmarking practices for multi-vector similarity search.
@@
Limitations. While we do not optimize for memory consumption of the multi-vector representations in this work, Lemur can be combined with scalar quantization to 8-bit precision and with methods reducing the number of stored multi-vector embeddings (e.g., Yan et al., 2025; Veneroso et al., 2025) that are orthogonal to it. We leave compatibility with extreme low-precision compression, such as 2-bit quantization, as future work.
@@ Acknowledgments
This work has been supported by the Research Council of Finland (grant #361902 and the Flagship programme: Finnish Center for Artificial Intelligence FCAI). The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.
@@ Appendix A LEMUR hyperparameters
In this section, we list the hyperparameters used for training the model used in Lemur (see Sec. 4). Lemur is robust with respect to the hyperparameters and we use the same hyperparameters for all experiments in this paper.
@@
Optimizer Adam (Kingma, 2014)
@@
Appendix B Ablation study additional results
B.1 Additional results for $k=100$
@@
In this section, we present additional results for the embedding dimension ablation study on all six BEIR datasets for $k=100$. For discussion, see Sec. 6.2.
B.2 Additional results for $k=10$
@@
In this section, we present additional results for the embedding dimension ablation study on all six BEIR datasets for $k=10$. The differences in recall between the configurations are similar as for $k=100$.
Appendix C End-to-end performance additional results
C.1 ColBERTv2
@@
In this section, we present the complete end-to-end performance results on all six BEIR datasets. On all six datasets, Lemur significantly outperforms the baseline methods. For further discussion, see Sec. 6.3.
C.2 Other multi-vector embedding models on Quora
@@
In this section, we perform the experiment in Sec. 6.3 using the four different modern multi-vector embedding models on the Quora dataset. The results are similar as for the SCIDOCS dataset.
Appendix D Effect of query distribution
D.1 Document encoder
@@
In this section, we repeat the end-to-end performance experiments of Sec. 6.3 with Lemur trained using a subset of the corpus $C_{1},\dots,C_{m}$ (encoded using the document encoder $D$) directly as a training set. This training method requires neither additional data, nor access to an encoder. Lemur still significantly outperforms the baseline methods.
D.2 Training set selection
@@
In this section, we study the effect of using a separate sample of actual queries for training Lemur. In the figures below, we consider Lemur trained using three different training sets (see Sec. 4.2):
@@
On both HotpotQA and ViDoRe, using a sample of actual queries yields a consistent performance improvement for Lemur.

### v2

@@ 7 Discussion
Because of their fine-grained representations that capture token-level interactions, multi-vector models are more expressive and, consequently, tend to have higher retrieval quality than single-vector models. However, the high latency of multi-vector retrieval engines has hindered their wider adoption. To address this issue, we introduce Lemur, a simple yet efficient framework that is an order of magnitude faster than earlier multi-vector similarity search methods, enabling efficient multi-vector retrieval at scale. Moreover, Lemur produces single-vector representations for queries and documents, reducing the task to single-vector similarity search. This makes it easy to integrate Lemur with state-of-the-art single-vector ANNS libraries.
@@
Because they are learned in a supervised fashion, the single-vector embeddings generated by Lemur are significantly more accurate than the fixed-dimensional encodings (FDEs) (Jayaram et al., 2024) that have been used for single-vector reduction earlier: in particular, $1024$-dimensional Lemur embeddings yield higher recall than $10240$-dimensional FDEs (see Section 6.2). This enables using lower-dimensional single-vector embeddings, which reduces the memory footprint and leads to lower end-to-end latency due to faster single-vector similarity search. Standard scalar- or product-quantization (Jégou et al., 2011) techniques can be applied to the single-vector embeddings generated by Lemur to further reduce memory consumption and end-to-end latency.
@@
The performance of earlier multi-vector similarity search methods has been evaluated almost exclusively on ColBERTv2 (Santhanam et al., 2022b) embeddings. We conduct a performance evaluation on four additional multi-vector text models and two multi-vector visual document retrieval models (see Section 6.3). Our results indicate that the model used to generate the embeddings has a larger impact on the performance of multi-vector similarity search methods than the embedded dataset. In particular, most earlier methods either fail to achieve acceptable recall levels at all or have prohibitively high latency on the LFM2-ColBERT-350M and answerai-colbert-small-v1 embeddings. In contrast, Lemur consistently achieves the highest recall levels with low latency regardless of the multi-vector model used to generate the embeddings. These observations underline the importance of establishing robust benchmarking practices for multi-vector similarity search.
@@
Limitations and future work. While we do not optimize for memory consumption of the multi-vector representations in this work, Lemur can be combined with e.g. product quantization (Jégou et al., 2011) or scalar quantization. We leave compatibility with extremely low-precision compression, such as $2$-bit quantization, as future work.
While the MLP in Lemur can be trained quickly and without a separate training set, even using the corpus itself as a training set may still pose an operational challenge when building an index incrementally as documents arrive. Our experiments show that using random hidden-layer weights is feasible (see Section 4.3), but we leave the investigation of synthetic and cross-corpus training sets as future work.
Finally, future work will examine whether the benefits of framing MaxSim as multi-output regression generalize to other similarity measures and retrieval tasks.
Impact Statement
This work improves the efficiency of multi-vector retrieval and is foundational machine learning research, and thus does not have a direct societal impact.
@@ Acknowledgments
This work has been supported by the Research Council of Finland (grant #361902 and the Flagship programme: Finnish Center for Artificial Intelligence FCAI). The authors acknowledge the research environment provided by ELLIS Institute Finland, as well as CSC – IT Center for Science, Finland, for computational resources.
@@ Appendix A LEMUR Hyperparameters
In this section, we list the hyperparameters used to train the Lemur model (see Section 4). Lemur is robust to the hyperparameters, and we use the same hyperparameters for all experiments in this paper. We did not observe benefits from using regularization because the model is a shallow network learning an exact, noise-free function.
@@
Optimizer Adam (Kingma and Ba, 2015)
@@
Appendix B LEMUR Universal Approximation
We assume that the MLP used in Lemur is universal on compact sets. Concretely, for any compact set $K\subset\mathbb{R}^{d}$, any continuous $h:K\to\mathbb{R}^{m}$, and any $\varepsilon>0$, there exists a hidden dimension $d^{\prime}$ such that
$\sup_{x\in K}\|h(x)-W\psi(x)\|_{\infty}\leq\varepsilon.$
This is the standard universal approximation property for MLPs on compact domains (Cybenko, 1989; Hornik, 1991).
Theorem B.1.
Let $K\subset\mathbb{R}^{d}$ be compact, and suppose every token-level query embedding lies in $K$. Then, for every $\varepsilon>0$, there exist a hidden dimension $d^{\prime}$, a feature map $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$, and a matrix $W\in\mathbb{R}^{m\times d^{\prime}}$ such that for all $x\in K$, $\|g(x)-W\psi(x)\|_{\infty}\leq\varepsilon$. Consequently, for every query $X\subset K$, $\|f(X)-W\Psi(X)\|_{\infty}\leq|X|\,\varepsilon$. Equivalently,
$\bigl|\mathrm{MaxSim}(X,C_{l})-\langle w_{l},\Psi(X)\rangle\bigr|\leq|X|\,\varepsilon$
for every document $l\in[m]$,
Proof.
For each fixed $l\in[m]$, the function $g_{l}$ is the maximum of finitely many linear functions, and hence it is continuous on $\mathbb{R}^{d}$. Therefore, $g:K\to\mathbb{R}^{m}$ is continuous on $K$.
By the universal approximation property, for any $\varepsilon>0$ there exist $d^{\prime}$, $\psi$, and $W$ such that
$\sup_{x\in K}\|g(x)-W\psi(x)\|_{\infty}\leq\varepsilon.$
Now fix any query $X\subset K$. We have
$f(X)-W\Psi(X)=\sum_{x\in X}\bigl(g(x)-W\psi(x)\bigr).$
Taking the $\ell_{\infty}$ norm and applying the triangle inequality yields
$\|f(X)-W\Psi(X)\|_{\infty}\leq\sum_{x\in X}\|g(x)-W\psi(x)\|_{\infty}\leq\sum_{x\in X}\varepsilon=|X|\,\varepsilon.$
Taking the $l$th coordinate gives
$\bigl|\mathrm{MaxSim}(X,C_{l})-\langle w_{l},\Psi(X)\rangle\bigr|\leq|X|\,\varepsilon.$
Appendix C Retrieval Quality Evaluation
In this section, we use nDCG@$10$ (Järvelin and Kekäläinen, 2002) to verify that Lemur preserves the retrieval quality of ColBERTv2 in the experiments of Section 6.3 and Appendix G.1. For each dataset, we list in the table below the nDCG@$10$ obtained using exact MaxSim search and then list the best QPS achieved by Lemur such that the retrieved results match the nDCG@$10$ obtained using exact search within the specified $\varepsilon$. Lemur is able to match the retrieval quality of ColBERTv2 while maintaining high QPS.
SCIDOCS FiQA TREC-COVID Touche Quora NQ HotpotQA
nDCG@$10$ 0.158 0.347 0.727 0.257 0.857 0.561 0.678
QPS ($\varepsilon<0.01$) 3142 2753 1994 2389 4068 917 108
QPS ($\varepsilon<0.001$) 1351 1038 708 2389 2191 179 77
Appendix D Correlation with Exact MaxSim
In this section, we measure the Pearson correlation coefficient (Pearson’s $r$) and the Spearman’s rank correlation coefficient by dataset (Spearman’s $\rho$) between the Lemur estimates and the exact MaxSim similarities, averaged over queries. Table 4 shows the correlation coefficients by dataset. The correlation coefficients are high ($\geq$ 0.942) on each dataset.
Table 4: Pearson correlation coefficient and Spearman’s rank correlation coefficient by dataset, averaged over queries.
Dataset Pearson’s $r$ Spearman’s $\rho$
HotpotQA 0.966 0.959
NQ 0.952 0.942
Quora 0.979 0.972
Touche-2020 0.963 0.951
TREC-COVID 0.994 0.990
FiQA 0.971 0.964
SCIDOCS 0.969 0.961
Appendix E Long Queries
While the default number of query vectors used in ColBERTv2 is 32, for long queries, e.g., in agentic applications, it can be preferable to use more vectors. To test the effectiveness of Lemur with longer queries, in the figure below, we use the ArguAna dataset from BEIR with 300 vectors per query, following Santhanam et al. (2022b). Lemur maintains its effectiveness in both retrieval performance and MaxSim approximation quality (Pearson correlation coefficient 0.987, Spearman’s rank correlation coefficient 0.983).
Appendix F Ablation Study Additional Results
F.1 Additional Results for $k=100$
@@
In this section, we present additional results for the embedding dimension ablation study on eight BEIR datasets for $k=100$. For discussion, see Section 6.2.
F.2 Additional Results for $k=10$
@@
In this section, we present additional results for the embedding dimension ablation study on eight BEIR datasets for $k=10$. The differences in recall between the configurations are similar to those for $k=100$.
Appendix G End-to-End Performance Additional Results
G.1 ColBERTv2
@@
In this section, we present the complete end-to-end performance results on eight BEIR datasets. On all eight datasets, Lemur significantly outperforms the baseline methods. For further discussion, see Section 6.3.
G.2 Other Multi-Vector Embedding Models on SCIDOCS
@@
In this section, we perform the experiment in Section 6.3 using the same four modern multi-vector embedding models on the SCIDOCS dataset. In addition, we also present results using the LateOn [^2] https://huggingface.co/lightonai/LateOn and mxbai-edge-colbert-v0-32m [^3] https://huggingface.co/mixedbread-ai/mxbai-edge-colbert-v0-32m models. The results are similar to those for the Quora dataset.
Appendix H Effect of Query Distribution
H.1 Document Encoder
@@
In this section, we repeat the end-to-end performance experiments of Section 6.3 with Lemur trained using a subset of the corpus $C_{1},\dots,C_{m}$ (encoded using the document encoder $D$) directly as a training set. This training method requires neither additional data nor access to an encoder. Lemur still significantly outperforms the baseline methods.
H.2 Training Set Selection
@@
In this section, we study the effect of using a separate sample of actual queries for training Lemur. In the figures below, we consider Lemur trained using three different training sets (see Section 4.3):
@@
On both HotpotQA and ViDoRe, using a sample of actual queries yields a consistent performance improvement for Lemur. However, Lemur is robust to the choice of training data.
H.3 Random Hidden Layer Weights
In this section, we perform an additional experiment in which we do not train the feature encoder $\psi$ at all. Instead, we use an MLP with random weights in the hidden layer, so the only component that adapts to the training data is the final projection layer. Specifically, we sample $R\in\mathbb{R}^{d^{\prime}\times d}$ with independent standard normal entries and use random features
$\psi_{\mathrm{ELM}}(x)=\sqrt{\frac{2}{d^{\prime}}}\,\mathrm{GELU}(Rx).$
This is often referred to as an extreme learning machine (ELM). The figure below shows the performance of the ELM version of Lemur compared to the version with a trained MLP on the NQ dataset. For discussion, we refer to Section 4.3.